OpenStudy (anonymous):

calc help, integral x/(sqrt(3-2x^(2)-x^(4))

mathslover (mathslover):

\[\large{\int{\frac{x}{\sqrt{3-2x^2-x^4}}}}\] Just for ease :)

OpenStudy (anonymous):

gracias lol

OpenStudy (anonymous):

The expression in the radical is simply a polynomial of x^2, so I'd let u=x^2.

OpenStudy (anonymous):

so far i got to \[\int\limits\limits_{n/a}^{n/a} x/\sqrt(4-(x ^{4}+1)^{2})\]

OpenStudy (anonymous):

thats (x^2+1)^2

OpenStudy (anonymous):

^

OpenStudy (anonymous):

its better to let \( u=1+x^2 \) i think :D

OpenStudy (anonymous):

came to 1/2 integral (du)/sqrt(4-(u)^2

OpenStudy (anonymous):

I would use arcsin integral correct?

OpenStudy (anonymous):

u/2 no ?! @mukushla

OpenStudy (anonymous):

sorry a little typo \(\large \frac{1}{2} \arcsin \frac{u}{2} \)

OpenStudy (anonymous):

and u r Quite Right @Ink412

OpenStudy (anonymous):

1/2 arcsin (u or 1+ (u)^2/ 2) + c?

OpenStudy (anonymous):

u mean 1+x^2 ?

OpenStudy (anonymous):

yeah lol

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

yw :)